OpenStudy (anonymous):

How do I determine if 7/16 is greater/less than 0.875?

OpenStudy (unklerhaukus):

to compare them you might want to get them (or some approximation) to the same form, [both as decimals for example] are there any fractions (in sixteenthths) that you can write as a decimal number?

OpenStudy (unklerhaukus):

the fraction i was thinking of is 8/16, can you convert that to a decimal?

OpenStudy (amistre64):

i was thinking cross multiplication :) 7/16 ??? 0.875 7 ??? 0.875(16)

OpenStudy (unklerhaukus):

how is it possible to multiply 0.875 by 16? @amistre64

OpenStudy (amistre64):

\[\frac{7}{16}~:~\frac{.875}{1}\] \[\cancel{16}\frac{7}{\cancel{16}}~:~16\frac{.875}{1}\] \[7~:~16(.875)\]

OpenStudy (amistre64):

since 50% of 16 = 8, is bigger than 7 ... then 87% of 16 is even bigger

OpenStudy (unklerhaukus):

but how can we simplify 6(.875) ?

OpenStudy (amistre64):

it just mutliplication .... learned it in the 3rd grade :) .875 16 ------ 5250 875 ------- 14.000

OpenStudy (amistre64):

maybe the 5th grade .... it was a long time ago for sure

OpenStudy (amistre64):

\[\frac{7}{16}\lt\left[\frac{1}{2}=.500\right]<.875\]

OpenStudy (anonymous):

7/16=0.4375

OpenStudy (unklerhaukus):